OpenStudy (anonymous):

Seven students applied for a position on the four-member student review board. In how many ways could four students have been selected for the board? please show me the steps.

OpenStudy (anonymous):

\[\dbinom{7}{4}\] \[\frac{7\times 6\times 5}{3\times 2\times 1}\] cancel first, multiply second

OpenStudy (anonymous):

is it 35

OpenStudy (anonymous):

you may notice that i chose to compute \[\dbinom{7}{3}\] which is the same thing because choosing 4 to serve is the same as choosing 3 to exclude

OpenStudy (anonymous):

yes, 35

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I have a question. why we have to exclude 4 in the denominator.Why can't be 7*6*5/4*3*2*1